OpenStudy (anonymous):

3x + y = 23 1/3 x - 1/4 y = 4 solve for x and y

hartnn (hartnn):

to eliminate the fraction, lets multiply 2nd equation by 12 what do u get?

OpenStudy (anonymous):

4 x - 3y = 16

hartnn (hartnn):

12*4=?

OpenStudy (anonymous):

48

hartnn (hartnn):

so it will be 4 x - 3y = 48 , isn't it ?

OpenStudy (anonymous):

oooh I see

hartnn (hartnn):

now to make the co-efficients of y equal in magnitude, multiply 1st equation by 3 and tell me what u get?

OpenStudy (anonymous):

9x + 3y = 69

hartnn (hartnn):

yes, now u have two equations : 9x + 3y = 69 4 x - 3y = 48 ADD them, and tell me what u get.

OpenStudy (anonymous):

13 x = 117 117/ 13 x= 9 I got it ! thanks a ton (:

hartnn (hartnn):

also find y welcome :)